Let π‘ˆ = {π‘Ž, 𝑒, 𝑖, π‘œ, 𝑒, 3, 6, 9, 12}, 𝐴 = {π‘œ, 𝑒, 3, 12}, 𝐡 = {𝑒, 9, 6, π‘Ž}, π‘Žπ‘›π‘‘ 𝐢 = {𝑒, 12, 𝑖, π‘œ}.


1. Find 𝐢 βˆͺ 𝐴


2. Find (𝐢′ ∩ 𝐡) βˆͺ 𝐢


3. Find {(𝐡′ βˆͺ 𝐴)β€² βˆ’ 𝐴 βˆͺ (𝐡 ∩ 𝐡′)}


1
Expert's answer
2022-04-08T07:04:13-0400

"1:\\\\C\\cup A=\\left\\{ o,u,i,3,12 \\right\\} \\\\2:\\\\\\left( C'\\cap B \\right) \\cup C=\\left( C'\\cup C \\right) \\cap \\left( B\\cup C \\right) =B\\cup C=\\left\\{ u,a,i,o,6,9,12 \\right\\} \\\\3:\\\\\\left( B'\\cup A \\right) '-A\\cup \\left( B\\cap B' \\right) =\\left[ \\begin{array}{c}\tB\\cap B'=F\\\\\tA\\cup \\left( B\\cap B' \\right) =A\\cup F=A\\\\\\end{array} \\right] =\\\\=\\left( B'\\cup A \\right) '-A=B''\\cap A'\\cap A'=B\\cap A'=B-A=\\left\\{ a,9,6 \\right\\} \\\\\\left[ those\\,\\,of\\,\\,B\\,\\,which\\,\\,are\\,\\,not\\,\\,in\\,\\,A \\right]"